A saturated solution of magnesium fluoride has a concentration [F]of 2.34 x 10-3 M. For this compound, Ksp = 1.17 x 10-3 O (1 please explain
Answer #1

Ans : Ksp = 4(1.17 x 10-3M)3​​​​​​​

MgF2 dissociates as :

MgF2 = Mg2+ + 2F-

If [F-] = 2.34 x 10-3 M

then [Mg2+] = 2.34 x 10-3 M / 2 = 1.17 x 10-3M

Ksp = [Mg2+] [F-]2

it is given as Ksp = 4S3

here S = 1.17 x 10-3M

so Ksp = 4(1.17 x 10-3M)3